OpenStudy (anonymous):

Compare and Contrast: Below are two equations. Solve each equation and compare the two solutions. Choose the statement that is true about each solution. Equation #1 Equation #2 8 − 5x = 4x − 1 4x + 4 = 10 − 2x The solution to equation #1 is smaller than the solution to equation #2. The solution to equation #1 is larger than the solution to equation #2. The solution to equation #1 is the same as the solution to equation #2. None of the statements above describes solutions to equations shown.

OpenStudy (anonymous):

rearrange the equations first 8 − 5x = 4x − 1 add 1 9-5x = 4x add 5x 9 = 9x divide by nine and you have the solution to equation 1 then rearrange equation 2 and compare them

OpenStudy (anonymous):

8 − 5x = 4x − 1 9x=9=> x=1 4x + 4 = 10 − 2x 6x=6=>x=1 The solution to equation #1 is the same as the solution to equation #2.

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

ur welcome again ;)